USING PYTHON. (Science: day of the week) Zeller’s congruence is an algorithm developed by Christian Zeller to calculate the day of the week. The formula is h = (q + 26(m+1)//10 + k + k//4 +j//4 +5j) % 7 where - h is the day of the week (0: Saturday, 1: Sunday, 2: Monday, 3: Tuesday, 4: Wednesday, 5: Thursday, 6: Friday). - q is the day of the month. - m is the month (3: March, 4: April, ..., 12: December). January and February are counted as months 13 and 14 of the previous year. - j is year//100. - k is the year of the century (i.e., year % 100). Write a program that prompts the user to enter a year, month, and day of the month, and then it displays the name of the day of the week.
Answer #1

#The formula is h = (q + 26(m+1)//10 + k + k//4 +j//4 +5j) % 7
#where - h is the day of the week
#(0: Saturday, 1: Sunday, 2: Monday, 3: Tuesday, 4: Wednesday, 5: Thursday, 6: Friday). - q is the day of the month.
#- m is the month (3: March, 4: April, ..., 12: December).
#January and February are counted as months 13 and 14 of the previous year.
#- j is year//100.
#- k is the year of the century (i.e., year % 100).

year = int(input('Enter the year: '))
month = int(input('Enter the month(1-12): '))
day = int(input('Enter the day of the month (1-31): '))

# As per the Zeller's congruence algorithm
#if January or February is entered we must add
# 12 to the month and minus 1 from the year. This
# will put us in month 13 or 14 of previous year

if month == 1:
month = month + 12
year = year - 1
elif month == 2:
month = month + 12
year = year - 1
  
#Lets calcualte the century values now
century = (year // 100)
century_year = (year % 100)

# Day of the week formula comes up to be
dotw = (day + ((26 * (month + 1)) // (10)) + century_year + ((century_year) // \
(4)) + ((century) // (4)) + (5 * century)) % 7

if dotw == 0:
print('Day of the week: Saturday')
elif dotw == 1:
print('Day of the week: Sunday')
elif dotw == 2:
print('Day of the week: Monday')
elif dotw == 3:
print('Day of the week: Tuesday')
elif dotw == 4:
print('Day of the week: Wednesday')
elif dotw == 5:
print('Day of the week: Thursday')
elif dotw == 6:
print('Day of the week: Friday')